97.5mL of gas is collected over water at 300C and 785mmHg. What is the pressure of the dry gas at 25.0C in a 105mL container?


1
Expert's answer
2021-05-03T07:39:27-0400

"\frac{P_1V_1}{T_1}= \frac{P_2V_2}{T_2}"


"\frac{785×97.5}{573}= \frac{P_2×105}{298}"


"P_2= 379.09 mmHg"